OpenStudy (anonymous):

Starting from rest, a 3-kg object reaches a speed of 8m/s in 2 seconds. What is the net force acting on the body ?(12 N in a direction of acceleration)

OpenStudy (shane_b):

\[V_f-V_i=at\]\[8m/s=a(2s)\]\[a=4m/s^2\]\[F=ma=3kg(4m/s^2)=12N\]So yes, I agree with your answer :)